OpenStudy (anonymous):

Please help with this? Solve the equation on the interval [0,2π) sin θ cos θ=-1/2

OpenStudy (anonymous):

So, for this you need to find the instances on the unit circle where Cos*Sin=-1/2

OpenStudy (anonymous):

So you need to look for where \[\sqrt{2}/2*-\sqrt{2}/2\] and vice versa. This locations are at 3pi/4 and 7pi/4

OpenStudy (anonymous):

Sooooo... i think you have two answers for this one.

OpenStudy (anonymous):

ok that works. the answers are 3pi/4 and 7pi/4 then right? the work made sense

OpenStudy (anonymous):

Yuppy

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

You're welcome! Again! haha